If sin t= 5/13, what is cos t?
Answer by jim_thompson5910(35256)

sin^2 + cos^2 = 1

(5/13)^2 + cos^2 = 1

25/169 + cos^2 = 1

cos^2 = 1 - 25/169

cos^2 = 144/169

cos = sqrt(144/169)

cos = 12/13

So cos(t) = 12/13